OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days. f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? (4 points)

OpenStudy (anonymous):

so is it 11.26 = 8(1.05)^n? and I work backwards to find n?

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

this is what you need to do: plug in numbers and see if its right: 11.26 = 8(1.05)^3 try this one and see...

undeadknight26 (undeadknight26):

then try this one: 11.26 = 8(1.05)^7 this is your answer.

OpenStudy (anonymous):

sorry undeadknight my wifi messed up.

OpenStudy (anonymous):

first one is 9.261

OpenStudy (anonymous):

second one is 11.25 Thankyou very much @undeadknight26

OpenStudy (anonymous):

can you answer another? @undeadknight26

OpenStudy (anonymous):

Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points)

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@HomeZone

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

Anyone?

undeadknight26 (undeadknight26):

Sorry i was afk. I think it is the amount grown.